OpenStudy (anonymous):

factor completly 49c^2+64-112c

OpenStudy (mertsj):

Rewrite it in descending order

OpenStudy (mertsj):

49c^2 - 112c + 64 There is no common factor so try to factor it as ( )( )

OpenStudy (anonymous):

so it would be then what i dont get itwould it be (49c^2)-(112c+64)

OpenStudy (mertsj):

When I see 49 I think of 7 times 7. And when I see 64 I think of 8 times 8. So try those and see if you get 112c for the middle term.

OpenStudy (anonymous):

(7c-8)(7c-8)

OpenStudy (mertsj):

Useless, quit trying to live up to your name. This is a factoring problem

OpenStudy (anonymous):

why dont i get it its probably right infront of me, you both are helping me

OpenStudy (mertsj):

It is very difficult to have two people trying to teach you something at the same time using two different methods. I will leave you to the capable hands of UselessMass.

OpenStudy (anonymous):

oh my god will one of you just help me with this problem and not fight

OpenStudy (anonymous):

Mertsj you said that when you see 49 you think of 7*7 which = 49 so i would write (49)(64)=112 am i somewhat on the right track

OpenStudy (anonymous):

thank you , i need to learn lol

OpenStudy (mertsj):

So do you understand now?

OpenStudy (anonymous):

so the formula is ax^2+bx+c=0 so in this case it would be 49^2+64=112 right?

OpenStudy (anonymous):

im sorry it would be 49^2+64x+112=0right

OpenStudy (anonymous):

u lost me

OpenStudy (anonymous):

yeah mertsj probably best you take over I don't have time to explain this I was hoping he had a grasp on the quadratic formula. Sorry for intruding.

OpenStudy (mertsj):

\[49c ^{2}-112c + 64\] is a trinomial. Trinomials come from multiplying two binomials. So \[49c ^{2}-112c + 64\]= (7c )(7c ) Both signs have to be negative since 112c is negative and the 64 if positive. Are you with me so far?

OpenStudy (anonymous):

im not a he , i kinda of do but im not good with it that much

OpenStudy (mertsj):

Now we have to find numbers whose product is 64. I think of 8 times 8. Do you?

OpenStudy (anonymous):

yeah so the begin is going to be (7c )(7c )

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

yes 8 and 8 is 64

OpenStudy (mertsj):

so let's try that. (7c-8)(7c-8). Now do you know what FOIL means?

OpenStudy (anonymous):

it would be 49c-64 right

OpenStudy (mertsj):

Do you know what FOIL means?????

OpenStudy (anonymous):

kinda let me be honest im not good at math sorry

OpenStudy (mertsj):

So just say no. It's ok.

OpenStudy (mertsj):

Could you multiply (x+3)(x-8)

OpenStudy (anonymous):

it wil come out to be a - right , if not then it will be 24

OpenStudy (mertsj):

So now I understand why you are having so much trouble with factoring. To multiply two binomials such as (x+3)(x-8) multiply the First term of each binomial. In this case it would be x times x which is x^2. That's where the F in FOIL comes from. Are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Now multiply the Outer terms which are x and -8. That gives you -8x. That is where the O of FOIL comes from. Still with me?

OpenStudy (mertsj):

Now multiply the Inner terms which are 3 and x. That gives you 3x. That is where the I of FOIL comes from. I will wait to see if you understand so far or not.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i have x-8 and x-3 right

OpenStudy (mertsj):

Now multiply the last terms of each binomial which are 3 and -8. That gives you -24.That is where the L of FOIL comes from.

OpenStudy (anonymous):

alright it would be -24

OpenStudy (mertsj):

So now we have x^2 -8x+3x-24 or if we add -8x + 3x we get x^2-5x -24.

OpenStudy (mertsj):

Yes. Very Good.

OpenStudy (mertsj):

So now we have a trinomial: x^2-5x -24. Suppose the directions said "Factor Completely." We would have to reverse the process of multiplication that we just did and write it as a product.

OpenStudy (anonymous):

huh

OpenStudy (mertsj):

(x+3)(x-8) = x^2 -5x -24 Are you still with me? If not, where did you get lost?

OpenStudy (anonymous):

i am but i will be right back i have to pick up someone up from work,dont go on yet

OpenStudy (mertsj):

Let me know when you are back.

OpenStudy (anonymous):

alright iht im back

OpenStudy (anonymous):

so where do i go from tthere